PLEASE HELP

Let G is a graph with 2n vertices and n^2 edges. An amicable pair of vertices is an unordered pair (u, v), such that dist(u, v) = 2. Prove that G has at least n(n − 1) amicable pairs of vertices.
Answer #1

Suppose v1,v2,v3.....vn are the n-vertices of graph.

We also know that maximum number of degree of any vertex is n-1.

d(v1)=(n-1)

d(v2)=(n-1)

...... d(vn)=(n-1)

Now ,i- 1 i- 1 We know thot of odd degmnees even Sum ol degree o a yertices in a qor aph G is cven eoual to e 1

Considered unordered pair(u,v) with dist(u,v)=2.

G has atleast n(n-1) amicable pair of vertices .

Complete graph is taken.